# ABG-NAS: Adaptive Bayesian Genetic Neural Architecture Search for Graph Representation Learning

### v1

@@ IV-A2 Parameter Settings
The ABG-NAS framework was configured with specific parameters to efficiently guide the GNN architecture search. For the expanded search space, the operation set $\mathcal{O}$ consisted of 5 P operations: GCN, SplineCNN, and three GraphSAGE variants (mean, max, and sum), as well as 8 T operations: ReLU, Linear, ELU, Sigmoid, Tanh, ReLU6, SoftPlus, and Leaky_ReLU. GCN was selected for its simplicity and efficiency in feature aggregation, while GraphSAGE and SplineCNN offer alternative aggregation and transformation mechanisms. For the AGOS algorithm, the depth $d$ of each individual in the population was set to range from 3 to 15. Both the population size $N_{s}$ and the maximum number of generations $N_{g}$ were set to 20. An early stopping strategy with a maximum of 300 epochs was used to train the GNN models for each individual. In BGTM, the interval $N_{b}$ was set to 5 generations. The hyperparameters optimized by BGTM and their search spaces were defined as follows: the hidden dimension size was set between $4$ and $256$, forward dropout rate was varied between $0.4$ and $0.6$, middle dropout rate was tuned between $0.2$ and $0.4$, the overall dropout rate was adjusted in the range of $0.3$ to $0.5$,the learning rate was adjusted in the range $10^{-4}$ to $10^{-1}$, and weight decay was tuned from $10^{-5}$ to $10^{-2}$.
IV-B Comparison with Experience-based GNN Architectures
@@
We first compare the performance of the AGOS-driven GNAS process with experience-based GNN architectures on three GNN models: GCN, SplineCNN, and GraphSAGE. The results are summarized in Table II. In the “Methods” column, EXP refers to a classical two-layer P-T-P-T architecture manually designed based on expert knowledge, where the P and T operations are derived from the base model shown in the first column. AGOS denotes the use of the proposed AGOS algorithm to automate the architecture search within the same operation set as the corresponding EXP baseline. The final row shows ABG-NAS, which operates on the expanded search space as defined in Subsection IV-A2.
@@
TABLE II: Comparison of accuracy (Acc) and macro F1 scores between experience-based (EXP) GNN architectures, AGOS-automated architectures, and ABG-NAS across four datasets. The base models include GCN, SplineCNN, and GraphSAGE. The table reports the mean accuracy and macro-F1 score over 10 independent runs, with standard deviations included.
Base Model Methods Cora[28] PubMed [29] CoraFull [30] Citeseer[29]
@@
The results in Table II demonstrate the clear advantages of the automated AGOS approach over manual GNN designs across all three base models and four datasets. For instance, on the Cora dataset, GCN’s accuracy improved from 86.76% to 87.93%, with the F1 score increasing from 85.65% to 86.40%. On PubMed, the accuracy rose from 86.57% to 89.05%, while the F1 score improved from 86.98% to 88.82%. These improvements indicate that the automated NAS process consistently discovers more effective architectures, adapting better to different datasets. ABG-NAS, leveraging a broader and more flexible search space, outperforms all models, achieving the best performance across all datasets, particularly excelling on larger datasets like CoraFull and Citeseer.
Impact of Dataset Characteristics. While the automated AGOS NAS method consistently outperforms the experience-based (EXP) methods across all datasets, the degree of improvement is closely related to the characteristics of the datasets themselves. As shown in Table I, the Cora dataset is the smallest, with only 2,708 nodes and 5,278 edges. The gains achieved by AGOS on Cora are relatively modest compared to larger and more complex datasets like PubMed and CoraFull. For instance, the accuracy improvement for GCN on Cora is 1.17% (87.93%-86.76%), whereas on PubMed and CoraFull, the improvements are 2.48% (89.05%-86.57%) and 1.59% (64.25%-62.66%), respectively. These substantial improvements on PubMed (19,717 nodes, 44,338 edges) and CoraFull (19,793 nodes, 126,842 edges) demonstrate that AGOS is particularly effective on larger, more complex graphs. This is because the flexibility in architectural design provided by NAS methods allows for the creation of sophisticated models that can effectively capture both local and global patterns.
@@
TABLE III: Comparison of ABG-NAS with seven popular GNAS methods (GraphNAS, SNAG, SANE, Auto-HeG, DFG-NAS, DE, and PSO) across four datasets. The table reports the mean accuracy and macro-F1 score over 10 independent runs, with standard deviations included. ABG-NAS consistently achieves the highest performance across all datasets, demonstrating its robustness and effectiveness. Results marked with * are official results reported in their respective papers.
Methods Cora[28] PubMed [29] CoraFull [30] Citeseer[29]
@@
GraphNAS* [8] 84.10±0.79 - 82.28±0.64 - - - 68.83±2.09 -
SNAG* [9] 81.01±1.31 - 83.24±0.84 - - - 70.14±2.40 -
SANE* [12] 84.25±1.82 - 87.82±0.57 - - - 74.33±1.54 -
Auto-HeG* [13] 86.88±1.10 - 89.29±0.27 - - - 75.81±1.52 -
DE [32] 86.0±0.01 84.0±0.03 88.0±0.00 88.0±0.00 66.0±0.00 52.0±0.00 75.0±0.00 72.0±0.00
PSO [33] 85.0±0.02 83.0±0.03 87.9±0.01 87.8±0.001 65.6±0.01 52.2±0.001 74.9±0.03 71.6±0.03
DFG-NAS[14] 86.6±0.03 84.6±0.05 88.9±0.01 88.7±0.01 70.4±0.03 60.9±0.05 77.2±0.05 73.4±0.06
ABG-NAS (Our) 88.32±0.02 87.52±0.05 89.88±0.04 89.67±0.02 72.14±0.06 63.03±0.07 77.44±0.02 73.50±0.01
IV-C Comparison with Diverse GNAS Methods
In this experiment, we compared the performance of ABG-NAS with seven state-of-the-art GNAS methods across four datasets. These methods cover a range of frameworks, including RL-based (e.g., GraphNAS[8]), gradient-based (e.g., SANE[12]), and EA-based approaches[32, 33, 14]. As shown in Table III, ABG-NAS consistently achieves superior results on all datasets. For instance, on the Cora dataset, ABG-NAS achieves an accuracy of 88.32%, outperforming the best baseline by 2.3%.
Although ACC is a fundamental metric for evaluating model performance, to more fairly demonstrate the advantages of ABG-NAS, we further introduce Macro-F1 as a complementary evaluation metric to compare ABG-NAS with other widely-used EA-based NAS methods.Specifically, on the Cora dataset, our method achieved a Macro-F1 score of 87.52%, significantly outperforming DFG-NAS (84.6%), DE (84.0%), and PSO (83.0%). Similarly, on PubMed, ABG-NAS achieved 89.67%, compared to 88.7% for DFG-NAS, 88.0% for DE, and 87.9% for PSO. Even on the larger and more complex CoraFull dataset, ABG-NAS maintained stability with a Macro-F1 score of 63.03%, outperforming methods that struggled on this dataset. Citeseer, with its high sparsity ($D_{\text{avg}}=2.84$) and 217 isolated nodes (as shown in Table I), poses a significant challenge for many NAS methods. Despite this, ABG-NAS achieves a Macro-F1 score of 73.50%, outperforming DFG-NAS (73.4%), DE (72.0%), and PSO (71.6%). This highlights ABG-NAS’s adaptability to sparse graphs, further showcasing its robustness across diverse graph structures.
@@ IV-C Comparison with Diverse GNAS Methods
In summary, the results from both Table III and Figures 5 and 4 underscore the advantages of ABG-NAS over state-of-the-art GA-based NAS methods (DFG-NAS), and other EA-based algorithms, i.e., DE, and PSO. Our method not only achieves higher accuracy and macro-F1 scores but also exhibits superior robustness and stability across diverse and complex graph structures. This makes ABG-NAS a more reliable choice for automated neural architecture search in GNNs, particularly for challenging real-world datasets.
TABLE IV: Comparison of ablation study results on three datasets. Each NAS framework was executed 10 times, and the average accuracy and Macro-F1 scores with standard deviations are reported. The full ABG-NAS framework includes CASS, AGOS, and BGTM, while the variants exclude one or more components to assess their impact. The best results are highlighted in bold.
Method Setting Cora[28] PubMed [29] CoraFull [30]
@@
Figure 6: Ablation study results showing F1 scores and percentage improvements for ABG-NAS variants across Cora, PubMed, and CoraFull datasets. Bar graphs illustrate F1 scores for each model variant, while line graphs depict the relative percentage improvements ($\Delta$F1) compared to the full ABG-NAS framework.
IV-D Ablation Study
@@
Compared to the baseline DFG-NAS, ABG-NAS introduces three key components: CASS, AGOS, and BGTM. CASS increases the flexibility and adaptability of GNN architectures across different graph structures, AGOS enhances the search process by balancing exploration and exploitation through more advanced genetic operations, and BGTM fine-tunes the hyperparameters to improve performance. In this subsection, we conduct an ablation study to validate the effectiveness of these components. We compare the full ABG-NAS framework with its three variants: ABG-NAS-V1 (DFG-NAS + CASS), ABG-NAS-V2 (DFG-NAS + AGOS), and ABG-NAS-V3 (DFG-NAS + CASS + AGOS). The baseline DFG-NAS [14] is used as the reference point. The results are summarized in Table IV.
@@
From Table IV, we observe that removing any component results in a consistent drop in Macro-F1 scores across all datasets, confirming the importance of each component in ABG-NAS. Notably, without the sophisticated AGOS search strategy, ABG-NAS-V1 (which only incorporates the expanded search space) performs worse than the baseline DFG-NAS across all datasets, indicating that CASS alone cannot drive improvements without an effective search strategy. In contrast, ABG-NAS-V2, which employs AGOS but lacks CASS, significantly outperforms DFG-NAS, highlighting AGOS’s capability to enhance performance even in a simpler search space. When CASS is added to ABG-NAS-V2, forming ABG-NAS-V3, there is a slight improvement over ABG-NAS-V2, demonstrating the benefit of combining CASS with a more advanced search algorithm. Finally, incorporating BGTM into ABG-NAS-V3 produces the full ABG-NAS framework, which delivers the best results overall, validating the critical role of BGTM in hyperparameter fine-tuning for optimal performance.
@@
We further visualize these results in Fig. 6, which highlights the F1 scores and the percentage improvements achieved by ABG-NAS over DFG-NAS and the three variants. While the performance gap between variants is smaller on simpler datasets like Cora and PubMed, it becomes more pronounced on complex datasets such as CoraFull. For instance, the full ABG-NAS framework achieves a 4.01% improvement in F1 score over DFG-NAS, a 7.84% improvement over ABG-NAS-V1, a 3.06% improvement over ABG-NAS-V2, and a 1.32% improvement over ABG-NAS-V3 on CoraFull. These results underscore that ABG-NAS, as a comprehensive and robust framework, is capable of addressing diverse graph structures and tasks, delivering superior performance across various datasets.
V Conclusion
@@
In this paper, we introduce ABG-NAS, a novel framework designed to efficiently optimize neural architecture search for improved graph representation learning. ABG-NAS seamlessly integrates three components: Comprehensive Architecture Search Space (CASS), Adaptive Genetic Optimization Strategy (AGOS), and Bayesian-Guided Tuning Module (BGTM). These components synergistically explore and exploit a diverse range of GNN architectures while jointly optimizing hyperparameters, effectively adapting to graph datasets with varying sparsity. Our framework is both model-agnostic and scalable, enabling it to support effectual downstream graph learning tasks across a variety of real-world datasets. Empirical evaluations on various benchmark datasets demonstrate that ABG-NAS consistently outperforms both manually designed GNN and existing state-of-the-art NAS methods. The results of our ablation study further validate the effectiveness of the composite design of our framework. In the future, we aim to further generalize the search space to better handle more complex and heterogeneous graph structures, while ensuring continued efficiency and scalability. Furthermore, we plan to develop an explainability layer to streamline the interpretability of learned representations, further enhancing the framework’s usability in practical applications.

### v3

@@ 4.1.2 Parameter Settings
The ABG-NAS framework was configured with specific parameters to efficiently guide the GNN architecture search. For the expanded search space, the operation set $\mathcal{O}$ consisted of 5 P operations: GCN, SplineCNN, and three GraphSAGE variants (mean, max, and sum), as well as 8 T operations: ReLU, Linear, ELU, Sigmoid, Tanh, ReLU6, SoftPlus, and Leaky_ReLU. GCN was selected for its simplicity and efficiency in feature aggregation, while GraphSAGE and SplineCNN offer alternative aggregation and transformation mechanisms.
For the AGOS algorithm, the depth $d$ of each individual in the population was set to range from 3 to 15. Both the population size $N_{s}$ and the maximum number of generations $N_{g}$ were set to 20. An early stopping strategy with a maximum of 300 epochs was used to train the GNN models for each individual.
To ensure fair and robust comparisons, all experiments were repeated across 10 independent runs using different random seeds. Additionally, no learning rate scheduler was applied for any method. A fixed learning rate was used throughout training to avoid scheduler-induced bias and maintain consistency across all baselines and the proposed framework.
In the Bayesian-Guided Tuning Module (BGTM), the update interval $N_{b}$ was empirically set to 5 generations, balancing tuning benefits and computational overhead.
The hyperparameters optimized by BGTM and their search spaces were defined as follows: the hidden dimension size was set between $4$ and $256$, forward dropout rate was varied between $0.4$ and $0.6$, middle dropout rate was tuned between $0.2$ and $0.4$, the overall dropout rate was adjusted in the range $0.3$ to $0.5$, the learning rate was adjusted in the range $10^{-4}$ to $10^{-1}$, and weight decay was tuned from $10^{-5}$ to $10^{-2}$.
In addition, the smoothing factor $\lambda=0.5$ used in the AGOS fitness control mechanism (Equation 10) was fixed to provide balanced influence between current and historical fitness trends. This value was empirically selected based on preliminary experiments for stability.
All methods were executed under the same computational budget, with 10 independent runs per dataset and consistent train/validation/test splits to ensure a fair comparison. The reported results are averaged over 10 runs, and standard deviations are included to indicate performance stability.
4.1.3 Compared Algorithms
Compared Algorithms. To ensure a fair and reproducible comparison, we provide below the configuration details of all baseline NAS methods evaluated in our experiments:
GraphNAS and SNAG: These reinforcement learning-based NAS methods were executed using their official implementations with default configurations. As these models are computationally intensive, we reused the results reported in their original papers (as marked with *).
SANE and Auto-HeG: These gradient-based methods were re-implemented using the authors’ official codebases. We retained all default settings without additional hyperparameter tuning.
DE: The differential evolution baseline was configured using standard defaults: scaling factor $F=0.8$ and crossover rate $CR=0.9$.
PSO: The inertia weight was fixed at $w=0.5$, while cognitive and social coefficients were set to $c_{1}=1.5$ and $c_{2}=1.5$, respectively. The velocity was bounded in [1,1][-1,1]. Coefficients $c_{1}$ and $c_{2}$ were selected via grid search within $[1.0,2.0]$.
DFG-NAS: The mutation probability was set to 0.3, and tournament selection size to 3, in accordance with the original implementation.
4.2 Comparison with Experience-based GNN Architectures
@@
We first compare the performance of the AGOS-driven GNAS process with experience-based GNN architectures on three GNN models: GCN, SplineCNN, and GraphSAGE. The results are summarized in Table 2. In the “Methods” column, EXP refers to a classical two-layer P-T-P-T architecture manually designed based on expert knowledge, where the P and T operations are derived from the base model shown in the first column. AGOS denotes the use of the proposed AGOS algorithm to automate the architecture search within the same operation set as the corresponding EXP baseline. The final row shows ABG-NAS, which operates on the expanded search space as defined in Subsection 4.1.2.
@@
Table 2: Comparison of accuracy (Acc) and macro F1 scores between experience-based (EXP) GNN architectures, AGOS-automated architectures, and ABG-NAS across four datasets. The base models include GCN, SplineCNN, and GraphSAGE. The table reports the mean accuracy and macro-F1 score over 10 independent runs, with standard deviations included.
[…]
@@
[…]
The results in Table 2 demonstrate the clear advantages of the automated AGOS approach over manual GNN designs across all three base models and four datasets. For instance, on the Cora dataset, GCN’s accuracy improved from 86.76% to 87.93%, with the F1 score increasing from 85.65% to 86.40%. On PubMed, the accuracy rose from 86.57% to 89.05%, while the F1 score improved from 86.98% to 88.82%. These improvements indicate that the automated NAS process consistently discovers more effective architectures, adapting better to different datasets. ABG-NAS, leveraging a broader and more flexible search space, outperforms all models, achieving the best performance across all datasets, particularly excelling on larger datasets like CoraFull and Citeseer.
Impact of Dataset Characteristics. While the automated AGOS NAS method consistently outperforms the experience-based (EXP) methods across all datasets, the degree of improvement is closely related to the characteristics of the datasets themselves. As shown in Table 1, the Cora dataset is the smallest, with only 2,708 nodes and 5,278 edges. The gains achieved by AGOS on Cora are relatively modest compared to larger and more complex datasets like PubMed and CoraFull. For instance, the accuracy improvement for GCN on Cora is 1.17% (87.93%-86.76%), whereas on PubMed and CoraFull, the improvements are 2.48% (89.05%-86.57%) and 1.59% (64.25%-62.66%), respectively. These substantial improvements on PubMed (19,717 nodes, 44,338 edges) and CoraFull (19,793 nodes, 126,842 edges) demonstrate that AGOS is particularly effective on larger, more complex graphs. This is because the flexibility in architectural design provided by NAS methods allows for the creation of sophisticated models that can effectively capture both local and global patterns.
@@
Table 3: Comparison of ABG-NAS with seven popular GNAS methods (GraphNAS, SNAG, SANE, Auto-HeG, DFG-NAS, DE, and PSO) across four datasets. The table reports the mean accuracy and macro-F1 score over 10 independent runs, with standard deviations included. ABG-NAS consistently achieves the highest performance across all datasets, demonstrating its robustness and effectiveness. Results marked with * are official results reported in their respective papers.
[…]
@@
[…]
4.3 Comparison with Diverse GNAS Methods
In this experiment, we compared the performance of ABG-NAS with seven state-of-the-art GNAS methods across four datasets. These methods cover a range of frameworks, including RL-based (e.g., GraphNAS[11]), gradient-based (e.g., SANE[17]), and EA-based approaches[38, 39, 19]. As shown in Table 3, ABG-NAS consistently achieves superior results on all datasets. For instance, on the Cora dataset, ABG-NAS achieves an accuracy of 88.32%, outperforming the best baseline by 2.3%.
Although ACC is a fundamental metric for evaluating model performance, to more fairly demonstrate the advantages of ABG-NAS, we further introduce Macro-F1 as a complementary evaluation metric to compare ABG-NAS with other widely-used EA-based NAS methods.Specifically, on the Cora dataset, our method achieved a Macro-F1 score of 87.52%, significantly outperforming DFG-NAS (84.6%), DE (84.0%), and PSO (83.0%). Similarly, on PubMed, ABG-NAS achieved 89.67%, compared to 88.7% for DFG-NAS, 88.0% for DE, and 87.9% for PSO. Even on the larger and more complex CoraFull dataset, ABG-NAS maintained stability with a Macro-F1 score of 63.03%, outperforming methods that struggled on this dataset. Citeseer, with its high sparsity ($D_{\text{avg}}=2.84$) and 217 isolated nodes (as shown in Table 1), poses a significant challenge for many NAS methods. Despite this, ABG-NAS achieves a Macro-F1 score of 73.50%, outperforming DFG-NAS (73.4%), DE (72.0%), and PSO (71.6%). This highlights ABG-NAS’s adaptability to sparse graphs, further showcasing its robustness across diverse graph structures.
@@ 4.3 Comparison with Diverse GNAS Methods
In summary, the results from both Table 3 and Figures 5 and 4 underscore the advantages of ABG-NAS over state-of-the-art GA-based NAS methods (DFG-NAS), and other EA-based algorithms, i.e., DE, and PSO. Our method not only achieves higher accuracy and macro-F1 scores but also exhibits superior robustness and stability across diverse and complex graph structures. This makes ABG-NAS a more reliable choice for automated neural architecture search in GNNs, particularly for challenging real-world datasets.
[…]
@@
[…]
Table 4: Comparison of ablation study results on three datasets. Each NAS framework was executed 10 times, and the average accuracy and Macro-F1 scores with standard deviations are reported. The full ABG-NAS framework includes CASS, AGOS, and BGTM, while the variants exclude one or more components to assess their impact. The best results are highlighted in bold.
[…]
4.4 Ablation Study
@@
Compared to the baseline DFG-NAS, ABG-NAS introduces three key components: CASS, AGOS, and BGTM. CASS increases the flexibility and adaptability of GNN architectures across different graph structures, AGOS enhances the search process by balancing exploration and exploitation through more advanced genetic operations, and BGTM fine-tunes the hyperparameters to improve performance. In this subsection, we conduct an ablation study to validate the effectiveness of these components. We compare the full ABG-NAS framework with its three variants: ABG-NAS-V1 (DFG-NAS + CASS), ABG-NAS-V2 (DFG-NAS + AGOS), and ABG-NAS-V3 (DFG-NAS + CASS + AGOS). The baseline DFG-NAS [19] is used as the reference point. The results are summarized in Table 4.
@@
From Table 4, we observe that removing any component results in a consistent drop in Macro-F1 scores across all datasets, confirming the importance of each component in ABG-NAS. Notably, without the sophisticated AGOS search strategy, ABG-NAS-V1 (which only incorporates the expanded search space) performs worse than the baseline DFG-NAS across all datasets, indicating that CASS alone cannot drive improvements without an effective search strategy. In contrast, ABG-NAS-V2, which employs AGOS but lacks CASS, significantly outperforms DFG-NAS, highlighting AGOS’s capability to enhance performance even in a simpler search space. When CASS is added to ABG-NAS-V2, forming ABG-NAS-V3, there is a slight improvement over ABG-NAS-V2, demonstrating the benefit of combining CASS with a more advanced search algorithm. Finally, incorporating BGTM into ABG-NAS-V3 produces the full ABG-NAS framework, which delivers the best results overall, validating the critical role of BGTM in hyperparameter fine-tuning for optimal performance.
@@
We further visualize these results in Figure 6, which highlights the F1 scores and the percentage improvements achieved by ABG-NAS over DFG-NAS and the three variants. While the performance gap between variants is smaller on simpler datasets like Cora and PubMed, it becomes more pronounced on complex datasets such as CoraFull. For instance, the full ABG-NAS framework achieves a 4.01% improvement in F1 score over DFG-NAS, a 7.84% improvement over ABG-NAS-V1, a 3.06% improvement over ABG-NAS-V2, and a 1.32% improvement over ABG-NAS-V3 on CoraFull. These results underscore that ABG-NAS, as a comprehensive and robust framework, is capable of addressing diverse graph structures and tasks, delivering superior performance across various datasets.
5 Conclusion
@@
In this paper, we introduce ABG-NAS, a novel framework designed to efficiently optimize neural architecture search for improved graph representation learning. ABG-NAS seamlessly integrates three components: Comprehensive Architecture Search Space (CASS), Adaptive Genetic Optimization Strategy (AGOS), and Bayesian-Guided Tuning Module (BGTM). These components synergistically explore and exploit a diverse range of GNN architectures while jointly optimizing hyperparameters, effectively adapting to graph datasets with varying sparsity. Our framework is both model-agnostic and scalable with respect to graph size and architectural complexity, enabling it to support downstream graph learning tasks across a variety of real-world datasets. Empirical evaluations on various benchmark datasets demonstrate that ABG-NAS consistently outperforms both manually designed GNN and existing state-of-the-art NAS methods. The results of our ablation study further validate the effectiveness of the composite design of our framework.
In the future, we aim to further generalize the search space to better handle more complex and heterogeneous graph structures, while ensuring continued efficiency and scalability across diverse graph sizes and model complexities. We also plan to conduct a comprehensive evaluation of ABG-NAS’s computational efficiency—such as GPU time and convergence speed—under unified benchmarking settings to strengthen the analysis of its practical viability. Furthermore, we plan to develop an explainability layer to streamline the interpretability of learned representations, further enhancing the framework’s usability in practical applications.
Acknowledgement
This work was supported in part by the Australian Research Council under Discovery Project Grant No. DP230100716.